OpenStudy (anonymous):

If f is the function whose graph is shown below, let h(x) = f(f(x)) and g(x) = f(x2). Use the graph of f to estimate the value of each derivative. find h'(2) and g'(2) i posted the graph in the comments

OpenStudy (anonymous):

heres the picture idk if it works

OpenStudy (michele_laino):

hint: we can write this: \[\Large \begin{gathered} h'\left( x \right) = f'\left( {f\left( x \right)} \right) \times f'\left( x \right) \hfill \\ \hfill \\ g'\left( x \right) = f'\left( {{x^2}} \right) \times \left( {2x} \right) \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

furthermore, from the graph, we get: \(f(2)=1\), and \(f(2^2)=f(4)=2\)

OpenStudy (anonymous):

i don't believe thats what i asked. i know how to derive h'(2) and g'(2) but i don't know how to find the values from the graph

OpenStudy (michele_laino):

the first derivative of f(x), is the slope of the tangent line to the graph at x=2

OpenStudy (anonymous):

yes i understand that but i don't know how to find those values. everything i try is wrong

OpenStudy (anonymous):

can you list the values for the slopes and i can take it from there.

OpenStudy (michele_laino):

as we can see, the slope of the tangent lines at x=1, and at x=2, are: \(4/5\) and \(1\) respectively

OpenStudy (michele_laino):

they are approximate value of course

OpenStudy (anonymous):

okay and what is the slope at x=4 (second part of the question)

OpenStudy (michele_laino):

I think it is \(4/2=2\) (approximated value)

OpenStudy (anonymous):

hi, both answers were wrong for part a and b is there any way you can re-calculate (i have one more attempt to answer them on webbing)

OpenStudy (michele_laino):

I don't know, since it is difficult to establish the slopes of tangent lines from the picture I suggest that you have to take a ruler, and then please draw the tangent lines you need, with the high accuracy, and measure the corresponding slopes

OpenStudy (anonymous):

is there any way the slopes can be calculate another way besides finding the points